# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v4

@@ Abstract.
We explain the structure of proof of our results on a problem that Sergio Doplicher posed in the 90s in the setting of conformal nets, and on a related problem posed by Yi-Zhi Huang in the setting of vertex operator algebras. The latter asks to find a direct construction of an equivalence due to Finkelberg between the quantum group modular fusion category at certain roots of unity and the fusion categories of affine vertex operator algebras at positive integer levels. The former asks for the construction of a quantum gauge group and field algebra for the category of localized endomorphisms of a conformal net, extending Doplicher-Roberts theory of the compact gauge group and field algebra in high dimensional algebraic QFT. In this summary we mostly focus on the affine VOA.
We solve Huang’s problem for all the Lie types for which it is known that the centralizer algebras of the non-negligible tensor powers of a given generating object $V$ for the quantum group fusion category (a list of $V$ for which fusion tensor powers are well defined has been given by Hans Wenzl for all the Lie types), is generated by the representation of the braid group associated to the $R$-matrix, and for Huang-Lepowsky ribbon braided tensor structure for the vertex operator algebras. In the classical type $A$, this property becomes the classical Schur-Weyl duality for the vector representation, and was used by Doplicher and Roberts to construct ${\rm SU}(d)$ as a special case of their unique compact gauge group and field algebra in algebraic QFT. Presently, known results in the literature on this generating property allow to construct our equivalence for the Lie types $A$, $B$, $C$, $D$, $G_{2}$. Our approach may be regarded in the setting of Connes noncommutative geometry.
We also solve a problem by Frenkel and Zhu on the construction of a (unitary) weak quasi-Hopf algebra structure on the Zhu algebra, associated to the minimum energy functor for all the Lie types, equivalent to that from quantum groups.
Our main tool is the construction of a unitary coboundary weak Hopf $C^{*}$-algebra in a sense extending Drinfeld coboundary quasi-Hopf algebras on a formal variable, associated to Wenzl functor for the quantum group fusion category, and of a Drinfeld twist. We apply a de-quantization continuous curve of Wenzl and a Drinfeld twist method to reach the Zhu algebra. The two Drinfeld coboundary symmetries we find in both contexts extend Doplicher-Roberts symmetric functors for rigid symmetric tensor categories with simple unit which play a central role in their duality theory.
@@ 1. Introduction
Moore and Seiberg [73], [74] discovered a connection between their equations and a certain structure of a tensor category arising from rational conformal field theories. The notion of ribbon braided tensor category has been settled in mathematics, and we refer to [23], [76], [30], [29], and, for their relations to compact quantum groups, to [80], and more generally to weak (quasi-) Hopf algebras to [13]. Turaev gave a definition of modular tensor category, based on previous work with Reshetikhin on the construction of knot and 3-manifold invariants see [96]. Several authors contributed to construct modular fusion categories form quantum groups at roots of unity, and a more complete historical account on the quantum group side may be found in [90], [92], see also the introduction of [13]. We denote by ${\mathcal{C}}({\mathfrak{g}},q)$ the resulting modular fusion category, with $q=e^{i\pi/d\ell}$, and $d\in\{1,2,3\}$ a suitable integer structure constant from Lie theory. Unitary tensor structures have been constructed by Wenzl and Xu [99], [105], and the former paper, as well as more recent papers by Wenzl and his coauthors play a major role in our work, and we try to explain this in this paper.
@@
On the side of CFT, our main interest in this paper is the Wess-Zumino-Witten model. Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians. A main difficulty in constructing tensor categories from models of conformal field theories is the construction of the associativity morphisms compatible with naturally assigned braiding morphisms via the two hexagonal equations. Another main difficulty is to show rigidity of the category.
@@
Let ${\mathfrak{g}}$ be a complex simple Lie algebra. These problems were first solved by Drinfeld in [25], [27] [28] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model. Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms. Furthermore he established rigidity of this category by showing a remarkable equivalence with the module category arising from the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$, which is a Hopf algebra. To construct the equivalence, he used Drinfeld coboundary matrix ${\overline{R}}$, which is a modification of the $R$-matrix by a square root of the coboundary defined by the ribbon structure.
@@
This is his remarkable Drinfeld-Kohno theorem, expanding an earlier theorem by Kohno [69] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types. Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ which is a Hopf algebra. As above mentioned, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
@@
Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras. The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals. In their work they consider both generic and non generic values of the levels. In particular they consider the negative integer levels $-k$, $k\in{\mathbb{N}}$, and proved a braided tensor equivalence of their category ${\mathcal{O}}_{-k}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the same type of roots of unity $q=e^{i\pi/d\ell}$, with the exception of some exceptional cases at low levels [66]. The order $\ell$ appearing in the root of unity and the positive integer $k$ are related by a well known equality, see e.g. Def. 20.1 and 21.2 in [13] and references therein.
@@
Shortly after, Finkelberg considered a certain semisimple subquotient $\tilde{\mathcal{O}}_{-k}$ of the representation category constructed by Kazhdan and Lusztig, paralleling the subquotient construction in the setting of quantum groups, and described a passage $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$ claimed as a braided tensor equivalence from the category $\tilde{\mathcal{O}}_{k}$ of certain integral modules of the affine Lie algebra at the positive integer level $k$ to the subquotient rigid braided tensor category $\tilde{\mathcal{O}}_{-k}$. Combining with the equivalence by Kazhdan and Lusztig, Finkelberg derived a connection between $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$. See the introduction and Theorem 4.3 in [33].
The original work by Finkelberg refers to the cited work in physics by Moore and Seiberg and in algebraic geometry by Beilinson, Feigin, and Mazur [4] for the rigid braided tensor structure on $\tilde{\mathcal{O}}_{k}$. As I learned from discussions with Huang, rigidity was not proved in [4], and there was no proof of rigidity using this approach.
By a private communication with Huang on the history of this problem, I learned that it finally turned out that the construction by Finkelberg described in [33] connecting $\tilde{\mathcal{O}}_{k}$ with Kazhdan and Lusztig structure inherited by $\tilde{\mathcal{O}}_{-k}$ could not suffice to derive rigidity of $\tilde{\mathcal{O}}_{k}$ from that of $\tilde{\mathcal{O}}_{-k}$. This implied that the proof of the equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$ was far from complete.
As later reported by Finkelberg in [34], it turned out that the completion of the proof of that proposition leading to rigidity of $\tilde{\mathcal{O}}_{k}$ and the equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$, involved the Verlinde formula as well (which became another main result in Finkelberg’s paper). This reflects also on Finkelberg composed equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}\to{\mathcal{C}}({\mathfrak{g}},q)$, with Kazhdan-Lusztig equivalence that moreover still passes through the negative integer levels. This implies in particular that the theorem does not include a few exceptional cases, including ${\mathfrak{g}}=E_{8}$ and $k=2$, which correspond to cases not treated in Kazhdan-Lusztig theory.
Following the introductions in [54], [61], Finkelberg work [33], [34] may be reinterpreted as giving the construction of a braided tensor category structure on $\tilde{\mathcal{O}}_{k}$. One may also conclude that these papers give a second proof of rigidity and modularity of this category.
On the other hand, Finkelberg equivalence theorem is stated as Theorem 7.0.2 in [1] without proof. The braided tensor equivalence between $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$ obtained from the combined works by Finkelberg, Kazhdan and Lusztig is considered indirect by several authors. This leads to the desire of constructing the equivalence in a natural, direct, and explicit way, and this problem is discussed by Y.-Z. Huang in [61] in great detail, where we select only a small part.
@@
We also mention Gannon’s observation in Sect. 6.2.3 of [38] regarding the desire of finding a direct connection between quantum groups and CFT.
From the side of mathematical axiomatizations of CFT, we work with respect to Huang and Lepowsky braided tensor structure for the representation category of vertex operator algebras.
@@
A direct construction of the equivalence may be helpful to see how rigidity or modularity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, or vice versa, by a given explicit equivalence. Or to simplify or give new construction of other structures, e.g. unitary structures, with a quantum group viewpoint, or to see more connections with quantum groups of models derived from or generalizing these fundamental models.
The study of how unitarization from the quantum group fusion categories [99] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence that passes through weak quasi-Hopf algebras was important for us since the early stages of applications of the theory of weak quasi-Hopf algebras [13], and further motivated us to study the above problem. Indeed one of our first results on unitarization was that weak quasi-Hopf algebras associated to fusion categories known to be equivalent allow to transport the unitary structure of one category to unitarize the other. In the process, knowledge of the explicit equivalence gives an understanding of the transported unitary structure. We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [13] for more details.
@@
Our starting point to approach this problem arose from a related problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory. We try to explain this problem.
Our approach to Huang problem is based on interpreting Finkelberg-Kazhdan-Lusztig equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the important class of WZW models that are described by affine Lie algebras at positive integer levels. We shall briefly come back to the previous history of these important models in physics.
@@
Algebraic quantum field theory had been initiated by Haag and Kastler in [47] in the 60s, and developed by founding breakthrough works by Doplicher, Haag, and Roberts [17], [18], [19], [20] in the 70s. We refer to the book by Haag [48], for an in-depth introduction to this theory, based on functional analysis, and more specifically on operator algebras, and also for more references.
@@
In this context, I learned from private discussions that for high dimensional quantum field theories categorical structures in their study first emerged from the discovery by Doplicher of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure. The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric strict tensor $C^{*}$-category with simple unit, starting from first principles in physics.
@@
Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality. In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to (space-like) infinity. This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map. From this the conjugate localized endomorphism $\overline{\rho}$ is derived. The left inverse may be described in the category as a map induced by the adjoint action of a solution of the conjugate equations on a intertwiner translated to the right by the tensor product. It was shown that the values of the statistics parameter (which is the constant $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), can be completely classified, and it can only take the values given by the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign which classifies charges of para-Bose or para-Fermi type. This analysis also needed to prove that the statistics parameter can not take the value 00 and was excluded by physically motivated conditions [20], [48].
@@
In a series of important papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality. Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid strict tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [21], [23],[22], and led to desired applications to high dimensional Algebraic QFT [24].

### v5

@@ Abstract.
This paper establishes connections between a problem posed by Doplicher and Roberts in the 1990s regarding construction of quantum gauge groups associated to the category of localized endomorphisms of conformal nets, and a problem by Huang concerning a direct proof of the Finkelberg equivalence theorem. We also unify the treatment of high dimensional algebraic QFT with the low dimensional case, resuming the study of Mack-Schomerus quantum symmetries, for which we previously proved a simplifying Hopf algebra property and studied the unitary structure of our global quantum gauge groups $A_{W}({\mathfrak{g}},q)$ of the WZW model. This program leads us to a solution of Huang’s problem in the setting of vertex operator algebras. Building upon our previous results, our approach utilizes a quantum gauge group framework to implement a unitary weak quasi-Hopf algebra structure on the Zhu algebra via a Drinfeld twist and Wenzl’s de-quantization continuous curve derived from the unitary coboundary weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ previously associated to the quantum group fusion category. Our construction is of an analytic nature, and was originally inspired by Drinfeld-Kohno theorem for Drinfeld category. Using this framework, in our previous work we established an independent rigid braided tensor structure on the module category of the affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ at positive integer level, providing a braided tensor equivalence with the modular fusion category of the corresponding quantum group at root of unity for all Lie types. In this paper, for the Lie types $A$, $B$, $C$, $D$, and $G_{2}$ we completely identify our structure with the Huang-Lepowsky braided tensor structure. The last step of the complete identification relies on the generating property of the braid group representation in the centralizer algebras of the tensor powers of the fundamental representation of the quantum group, which is currently known for these stated Lie types. This reduction builds on uniqueness results of braided tensor structures in presence of a generating object.
@@ 1. Introduction
1.1. Historical Background: Conformal Field Theories and Quantum Groups
Primary fields, introduced under a different name by Mack and Salam in the context of higher-dimensional conformal field theories [68], were investigated in 2D by Belavin, Polyakov, and Zamolodchikov. They recognized that the local symmetry algebra is given by the infinite-dimensional Virasoro algebra [5], building upon previous work by Feigin and Fuchs [29] and Kac [62] on its representation theory.
Witten introduced an interacting $2D$ conformal field theory, now known as the Wess-Zumino-Witten (WZW) model [103]. He showed that this model possesses a symmetry satisfying the relations of an affine Kac-Moody algebra. In the same year, Knizhnik and Zamolodchikov showed that the correlation functions of a WZW model satisfy a system of partial differential equations [65]. This Kac-Moody framework was related to the Virasoro algebras by Goddard, Kent and Olive [37], who introduced the coset construction to realize Virasoro minimal models directly from WZW models.
The mathematical formulation of the KZ equations was made rigorous by Tsuchiya and Kanie [95], who studied the model for ${\hat{\mathfrak{s}l}}_{2}$ and showed that the associated monodromy yields representations of the braid group that factor through the Hecke algebra at a root of unity.
The discovery of quantum groups first arose in physics literature in the work by Faddeev, introduced as noncommutative function algebras. In the mid 80s breakthrough papers by Drinfeld and Jimbo discovered them as quantization of universal enveloping algebras, and developed their properties. We refer to [8] for an in-depth presentation and historical information.
Kohno proved that the monodromy representations of the KZ equations are equivalent to those derived from the universal $R$-matrix of a quantum group [66]. Before recalling the subsequent deep developments of the theorem by Kohno, we need first review the emergence of tensor categories in constructive conformal field theory.
Moore and Seiberg [73], [74] derived certain equations from rational conformal field theory and discovered a connection between their equations and the structure of a tensor category. The corresponding mathematical framework, specifically the notion of ribbon braided tensor category, has since been settled in mathematics, we refer to [26], [27], and references therein.
An independent derivation of strict braided tensor $C^{*}$-categories arose from the axiomatic operator algebraic approach to QFT in the 1970s, see [20], [76], [45] and references therein. The connections between tensor $C^{*}$-categories and compact quantum groups are studied [80].
@@ 1.1. Historical Background: Conformal Field Theories and Quantum Groups
Turaev gave a definition of modular tensor category, based on previous work with Reshetikhin on the construction of knot and 3-manifold invariants see [96]. Several authors contributed to construct modular fusion categories form quantum groups at roots of unity, and a more complete historical account on the quantum group side may be found in [90], [92], see also the introduction of [10]. We denote by ${\mathcal{C}}({\mathfrak{g}},q)$ the resulting modular fusion category, with $q=e^{i\pi/d\ell}$, and $d\in\{1,2,3\}$ a suitable integer structure constant from Lie theory. Unitary tensor structures have been constructed by Wenzl and Xu [99], [105]. Wenzl’s work plays a major role in our work, and we try to explain this in this paper.
@@
In the constructive approach to CFT, the Wess-Zumino-Witten model is of central interest. Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians. A main difficulty in constructing tensor categories from models of conformal field theories is the construction of the associativity morphisms compatible with naturally assigned braiding morphisms via the two hexagonal equations. Another main difficulty is to show rigidity of the category.
@@
Non-strict examples motivated by conformal field theory arose from the work of Drinfeld on quasi-Hopf algebras [24]. Let $\mathfrak{g}$} be a complex simple Lie algebra. The mentioned problems were first solved by Drinfeld in [22], [24] [25] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the KZ differential equation. Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms. Furthermore he established rigidity of this category by showing a remarkable ribbon, braided tensor equivalence with the module category arising from the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$, which is a Hopf algebra. To construct the ribbon, braided, tensor equivalence, he used Drinfeld coboundary matrix $\overline{R}$}, which is a modification of the $R$-matrix by a square root of the coboundary defined by the ribbon structure.
@@
This is his remarkable Drinfeld-Kohno theorem, expanding the earlier theorem by Kohno [66]. Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$. In particular, the representations of the braid group arising from monodromy of the KZ equation are shown equivalent to those arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types.
@@
Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras. The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to $\mathfrak{g}$} at shifted levels excluding the non negative rationals. For convenience we denote by -$\ell$ the shifted levels allowed by Kazhdan and Lusztig. In their work they proved a braided tensor equivalence of their category $\mathcal{O}$$_{-\ell}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity $q=e^{i\pi/d\ell}$, with $d$ the ratio of the square lenghts of the long to the short root, with the exception of some cases [63].
@@
With $k$ a positive integer, defined as the physical level, and $\ell=k+\check{h}$, with $\check{h}$ the dual Coxeter number of $\mathfrak{g}$}, Finkelberg considered a certain semisimple subquotient $\tilde{\mathcal{O}}_{-\ell}$ of the representation category constructed by Kazhdan and Lusztig, paralleling the subquotient construction of the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ in the setting of quantum groups at roots of unity. He defined an equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}$ from the category $\tilde{\mathcal{O}}_{\ell}$ of certain integral modules of the affine Lie algebra at level $k$ introduced by Beilinson, Feigin, and Mazur to the subquotient rigid braided tensor category $\tilde{\mathcal{O}}_{-\ell}$. Combining with the equivalence by Kazhdan and Lusztig, Finkelberg derived an equivalence between $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$. See the introduction and Theorem 4.3 in [30].
While the equivalence between quantum group fusion categories and categories of modules of affine Lie algebras at positive integer level was proposed in the mid 1990s, the complete proof of rigidity of $\tilde{\mathcal{O}}_{\ell}$ and derivation of the tensor equivalence was given in 2013 using the Verlinde formula [31] building on work by Faltings and Teleman. We refer the reader to the appendix for more historical remarks concerning the proof of rigidity for affine vertex operator algebras and affine Lie algebras at positive integer levels.
The ribbon braided tensor equivalence established in [30], [31] relies on a passage through negative levels and then applies the equivalence with the quantum group fusion category by Kazhdan and Lusztig. As a consequence, Finkelberg’s proof excludes certain exceptional cases like $\mathfrak{g}=E_{8}$ at level $k=2$ since this case is excluded in [63].
1.2. Huang’s Problem
This led us to consider the following problem posed by Y.-Z. Huang in [58].
@@ 1.2. Huang’s Problem
Moreover, in light of Drinfeld’s construction of the ribbon braided tensor equivalence in Drinfeld-Kohno theorem, and his consequent derivation of rigidity of the Drinfeld category, is natural to ask whether the Verlinde formula is necessary to prove rigidity, and whether rigidity is necessary to prove the ribbon braided tensor equivalence $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$.
@@
We also mention Gannon’s observation in Sect. 6.2.3 of [35] regarding the desire of finding a direct connection between quantum groups and CFT. Our work provides a direct constructive method–addressing Huang’s Problem 4.4 in [58]–by utilizing a global quantum gauge group method originating in algebraic quantum field theory, and the Zhu algebra to establish the equivalence.
From the side of mathematical axiomatizations of CFT, Huang and Lepowsky have done extensive work on the construction of tensor structures of modules and showing all the properties of modular tensor category [47], [48], [49], [50], [51], [52], [53], [54], [55], [56].
@@
An explicit construction of the equivalence is helpful to see how rigidity, modularity or unitarity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, by our explicit equivalence.
1.3. Inspiration from Algebraic Quantum Field Theory
@@
Our starting point to approach this problem arose from a related problem posed by Doplicher and Roberts in the axiomatic setting of the operator algebraic approach to quantum field theory (AQFT).
Problem posed in Section 7 in [21]. Can we extend the duality theory for compact groups [61] Theorem 6.1 to more general objects, by replacing strict symmetric monoidal C*-categories by strict braided monoidal C*-categories? What is the full class of compact group-like objects which arise in this way?
@@
We next explain this problem and connect it to Huang’s problem. Algebraic quantum field theory was initiated by Haag and Kastler in [44] in the 1960s, and developed by founding breakthrough works by Doplicher, Haag, and Roberts [14], [15], [16], [17] in the 1970s. The operator algebraic approach to QFT joins high rigour provided by the operator algebra methods with first principles in physics on which the theory is formulated. We refer to the book by Haag [45], for an in-depth introduction to this theory.
@@
In the context of high dimensional quantum field theories, categorical structures first emerged from the discovery of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure. The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric strict tensor $C^{*}$-category with simple unit.
@@
Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality. In their work rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to (space-like) infinity. This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map. From this the conjugate localized endomorphism $\overline{\rho}$ is derived. The left inverse may be described in the category as a map induced by the adjoint action of a solution of the conjugate equations on a intertwiner translated to the right by the tensor product. It was shown that the values of the statistics parameter (which is the constant $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), can be completely classified, and it can only take the values given by the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign which classifies charges of para-Bose or para-Fermi type. This analysis also needed to prove that the statistics parameter can not take the value 0 and was excluded by physically motivated conditions [17], [45].
@@
In a series of papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality. Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid strict tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [18], [20],[19], and led to desired applications to high dimensional Algebraic QFT [21].
@@
1.4. Quantum symmetries and primary field reconstruction in low-dimensional AQFT